A furniture store marks up the price that it pays the manufacturer for dining room sets by 30%. If the furniture store puts a selling price of $2499.99 on a dining room , what did the furniture store pay for the set? Round your answer to the nearest cent.
Answer by richwmiller(17219)

1.3x=2499.99
x=2499.99/1.3=1923.069
rounded 1923.07
check
30% of 1923.07=576.92
1923.07+576.92=2499.99
ok